4) Find each of the two areas bounded by the curves y=x³-4x and y = x² + 21.

1
Expert's answer
2022-05-30T05:37:49-0400

Find each of the two areas bounded by the curves "y=x^3-4x" and"y=x^2+2x."



"x^3-4x=x^2+2x"

"x(x^2-x-6)=0"

"x(x+2)(x-3)=0"

"x_1=-2, x_2=0, x_3=3"

"A_1=\displaystyle\int_{-2}^{0}(x^3-4x-(x^2+2x))dx"

"=[\dfrac{x^4}{4}-\dfrac{x^3}{3}-3x^2]\begin{matrix}
 0 \\
 -2
\end{matrix}"

"=0-(\dfrac{(-2)^4}{4}-\dfrac{(-2)^3}{3}-3(-2)^2)"

"=\dfrac{16}{3}({units}^2)"


"A_2=\displaystyle\int_{0}^{3}(x^2+2x-(x^3-4x))dx"

"=[-\dfrac{x^4}{4}+\dfrac{x^3}{3}+3x^2]\begin{matrix}
 3 \\
 0
\end{matrix}"

"=-\dfrac{(3)^4}{4}+\dfrac{(3)^3}{3}+3(3)^2-0"

"=\dfrac{63}{4}({units}^2)"